OpenStudy (anonymous):

Anita can paint the shed in 5 hours. If Kris helps her, they can paint the shed in 3 hours. How long would it take Kris to paint the shed alone?

OpenStudy (nikvist):

\[\frac{1}{T}=\frac{1}{3}-\frac{1}{5}=\frac{2}{15}\quad\Rightarrow\quad T=\frac{15}{2}=7.5\,h\]